Where do you put the parentheses to make this true?
5-3x2+8x5-2;28

Answer by stanbon(57377)

Where do you put the parentheses to make this true?
5-3x2+8x5-2;28
(5-3)x2+8x(5-2)=28
2x2+8x3=28
4+24=28
28=28
Cheers,
Stan H.